OpenStudy (anonymous):

prove: sin Ɵ-sin Ɵ-cos^2 Ɵ=sin^3Ɵ someone help please I am so stuck

OpenStudy (solomonzelman):

The way you wrote the question.... I am not going to assume or redetermine anything. \(\Large\color{black}{ \bf sin Ɵ-sin Ɵ-cos^2 Ɵ=sin^3Ɵ }\) \(\Large\color{black}{ \bf -cos^2 Ɵ=sin^3Ɵ }\) \(\Large\color{black}{ \bf -(1-sin^2 Ɵ)=sin^3Ɵ }\) \(\Large\color{black}{ \bf -1+sin^2 Ɵ=sin^3Ɵ }\) \(\Large\color{black}{ \bf sin^3Ɵ +1-sin^2 Ɵ=0 }\) \(\Large\color{black}{ \bf sin^3Ɵ -sin^2 Ɵ+1=0 }\) say, let sinƟ=a and solve. Let me know if you need more help.

OpenStudy (anonymous):

I dont understand this at all would that be the answer?0.o

OpenStudy (solomonzelman):

\(\Large\color{black}{ \bf a^3 -a^2 +1=0 }\)

OpenStudy (anonymous):

you cant subtract the a^3-a^2 right? so do i just move the 1 over?

OpenStudy (solomonzelman):

use Newton's Method